A population of 100 crayfishes were observed in the wild. Three colors exist in this population; 47 red (RR), 13 orange (RY), and 40 yellow (YY).

A. Calculate the observed total number of each allele in the population (1 point) and the observed allele frequencies (1 points).

B. Using the Hardy-Weinberg equilibrium equation (p2 +2pq+q2=1.0) calculate what the predicted genotype frequencies would be if under Hardy-Weinberg equilibrium (2 points) and calculate the predicted number individuals of each genotype (2 points).

C. How do the observed genotypes of the population differ from the predicted genotypes (2 points)?

D. What conclusions can you make based on the differences you see in your observed population (2 points)?
Answer #1

RR = 47. RY = 13 YY = 40 100 107 A. # R = (49X4)+ 13 ee #Y= (40x+)+13 93 S 93 0.465 ofR = 107 - 0-535 y = doo exp GF # Bo Pre

Please rate high.